OpenStudy (anonymous):

Griselda budgets $50 per month before sales tax for clothes. How much will she spend on clothes each month if she is charged 6% sales tax

OpenStudy (anonymous):

50 +50*0,06=53

OpenStudy (anonymous):

Keisha's parents want to save $25,000 in her college savings account over the next 20 years. They have $10,000 to use as an initial deposit. What simple annual interest rate do they need to meet their goal? Round your answer to the nearest tenth.

OpenStudy (anonymous):

a)4.5 b)5.0 c)7.5 d)12.5

OpenStudy (anonymous):

10000 + 10000*20r=25000 200000r=15000 r=15/200 =0,075 7,5%

OpenStudy (anonymous):

Would you make more interest by investing $3000 in a 6-month certificate of deposit (CD) at 3% interest or by investing $1000 in a 12 month CD at 3.5% interest? Show your work.